OpenStudy (fanduekisses):

CHECK MY WORK PLEASE? Calculus Rate of Change:

OpenStudy (fanduekisses):

Indy is 6 feet tall and is walking at a rate of 3 feet per second towards a lamppost that is 18 feet tall. At what rate is his shadow due to the lamppost shortening when he is 6 feet from the base of the lamppost?

OpenStudy (fanduekisses):

I got -6 ft/sec

OpenStudy (fanduekisses):

\[\frac{ 18 }{ w+s }=\frac{ 6 }{ w }\]

OpenStudy (fanduekisses):

\[12\frac{ dw }{ dt }=6\frac{ ds }{ dt }\]

OpenStudy (fanduekisses):

\[\frac{ ds }{ dt }= -6ft/\sec\]

OpenStudy (anonymous):

what grade r u in

OpenStudy (anonymous):

a higher grade than i am XP

OpenStudy (anonymous):

i am in 7

OpenStudy (anonymous):

same

OpenStudy (anonymous):

lol @shyguy1026 will you be my fan please

OpenStudy (anonymous):

ssssssooooooorrrrrrryyyyyy

OpenStudy (m3ga-bit3):

@kisses what u tryin to figure out?

OpenStudy (anonymous):

I had asked you what grade you were dont be so mean man

OpenStudy (anonymous):

sorry i would help but cant :C

OpenStudy (fanduekisses):

I'm trying to find dh/dt

OpenStudy (m3ga-bit3):

oookkk lemme see...

OpenStudy (fanduekisses):

Ooops sorry not that lol

OpenStudy (fanduekisses):

****I'm looking for the rate of change in which the shadow is shortening.

OpenStudy (fanduekisses):

Similar triangles

OpenStudy (anonymous):

sorry i have to do my school work see youll in a bit :P

OpenStudy (m3ga-bit3):

ummm....do have the triangles so i can see them?

OpenStudy (fanduekisses):

I found this online http://img.sparknotes.com/figures/6/6feb2790f92651105afa8c668bf0b7ba/figure3-2p-3.gif

OpenStudy (fanduekisses):

But I used different variables s for x(t) and w for y(t)

OpenStudy (m3ga-bit3):

woow...um lemme see if i can find one of my ppl to help u cuz i dont wanna do it wrong

OpenStudy (fanduekisses):

ok

OpenStudy (anonymous):

That's what I got. I may be wrong though let me double check

OpenStudy (matlee):

yea the andwer is right

OpenStudy (xapproachesinfinity):

 try to picture things the shadow is changing with time when Indy gets closer and closer to lamppost

OpenStudy (xapproachesinfinity):

Oh you get the picture above! now you need to connect the dots what do you think you need to connect x(t) with y(t)

OpenStudy (xapproachesinfinity):

Hint: use triangle similarity

OpenStudy (fanduekisses):

Yes I used triangle similarity, ratios

OpenStudy (fanduekisses):

\[\frac{ x(t)+y(t) }{ 18 }=\frac{ x(t) }{ 6 }\]

OpenStudy (fanduekisses):

so \[ 6\frac{ dy(t) }{ dt }=12\frac{ dx(t) }{ dt }\]

OpenStudy (xapproachesinfinity):

ok good but remember that dx/dt is negative

OpenStudy (xapproachesinfinity):

decreasing means you are having a negative slope

OpenStudy (fanduekisses):

\[\frac{ dx(t) }{ dt }= -2 ft/\sec\]

OpenStudy (fanduekisses):

ooops I forgOt tO plUg IN LOL

OpenStudy (xapproachesinfinity):

i didn't do the calculations so I'm not sure what it is but as long as you did those ratios step the right way and put the right givens then you should be good. I'm out of time so i won't be able to check your math to see it is correct good luck

OpenStudy (xapproachesinfinity):

yes you need to plug in the givens

OpenStudy (fanduekisses):

So yea the rate of change of the shadow is -6 ft/sec

OpenStudy (fanduekisses):

Ok Thank you :D

OpenStudy (xapproachesinfinity):

Yes it should be sloping down so negative is good

OpenStudy (fanduekisses):

Have a nice day!

OpenStudy (xapproachesinfinity):

You are welcome and thank you!